### v1

@@ 1 Introduction
Medical image segmentation (MIS) has great practical value as it can assist relevant medical staff in locating the lesion area and improve the efficiency of clinical treatment. In recent years, UNet (Ronneberger et al, 2015), an encoder-decoder model based on U-shape architecture, has been extensively employed for MIS due to its strong scalability. As a result, many studies have been conducted based on the U-shape architecture, such as UNet++ (Zhou et al, 2018), which reduces the semantic gap between the encoder and decoder by incorporating dense connections. Additionally, Att-UNet (Oktay et al, 2018) introduces a gating mechanism to enable the model to focus on specific targets.
@@
The above improvements are all based on CNNs, which suffer from poor global information acquisition due to the natural locality of the convolution operation. In contrast, ViT (Dosovitskiy et al, 2020a) incorporates a self-attention mechanism (SA) that enhances the modeling ability of long-range dependencies and effectively captures holistic image semantic information, making it an ideal choice for intensive prediction tasks like image segmentation. As a result, recent advances can be categorized into two types. On the one hand, hybrid structures that combine CNNs and ViTs have gained popularity. For example, UCTransNet (Wang et al, 2022a) replaces the skip connection in UNet with the CTrans module, alleviating the issue of incompatible features between the encoder and decoder. MT-UNet (Wang et al, 2022c) employs CNNs at a shallow level and Local-Global SA along with the external attention mechanism at a deep level, to obtain richer representation information. On the other hand, some studies have leveraged pure ViTs, such as Swin-UNet (Cao et al, 2021), which replaces the convolution operation in U-Net with the Swin Transformer Block and achieves superior results.
@@
Although the aforementioned models have achieved notable successes in medical image area or other vision area (Xiang et al, 2022, 2023b), they mainly focus on the spatial domain, and there is a lack of research on medical image segmentation (MIS) from the frequency domain perspective. In general computer vision, feature extraction in the frequency domain has been shown to be a powerful approach (Zhong et al, 2022; Rao et al, 2021). For instance, GFNet (Rao et al, 2021) leverages 2D discrete Fourier transform (2D DFT) to convert features from the spatial domain to the frequency domain, and filters in the frequency domain are used for learning representation. For MIS, many pathological areas are difficult to extract in the spatial domain, while they can be more easily extracted in the frequency domain (Huang et al, 2021). However, previous methods extract the frequency domain information only in a single axis, resulting in some signals that are still difficult to distinguish, which can be explained in Figure 1. In Figure 1 (a), it is challenging to visually distinguish the three regions from the spatial domain perspective. In contrast, Figure 1 (b) illustrates the signal strength of the three regions in the frequency domain, revealing relatively apparent differences in signal strength among them. However, there are still some signal strength intersections between the green and red curves in the case of employing DFT only for a single axis (the Height-Width axis). By considering the frequency domain signal strength of the three axes together, as shown in Figure 1 (c), it becomes evident that there is no signal strength intersection among the three curves. Motivated by this observation, we propose to extract and fuse features using a multi-axis approach.
@@
    The Multi-Axis External Weights block is proposed to obtain global and local information simultaneously, and frequency domain feature signals are introduced to comprehend the context more effectively.
@@
The remainder of this paper is structured as follows. In Section 2, we give the related works related to the medical image segmentation and ViT based techniques, then briefly introduce our method. In Section 3, The details of our multi-axis external weights UNet based on the U-shape architecture are presented. Extensive evaluations compared with state-of-the-art methods and comprehensive analyses of the proposed approach are reported in Section 4. Finally, the conclusion of this paper and discussion of future works are presented in Section 5.
@@ 2.1 Medical Image Segmentation
First, a lot of endeavours have been concentrating on the segmentation process. These methods conquer different restrictions on conventional medical division techniques. Yet, there isn’t so much as one single strategy to be considered a better technique for different kinds of images; those techniques are only suitable for particular images and other applications (Xiang et al, 2023d).
@@ 2.2 ViT Based Techniques
Transformer models (Dosovitskiy et al, 2020b) entirely rely on the self-attention mechanism to build long-distance dependencies, which have achieved great success in almost all natural language processing tasks (Gao et al, 2023). Vision transformer (ViT) (Xiang et al, 2023a) is one of the earlier attempts to introduce transformer models into vision tasks, which applies a pure transformer architecture on non-overlapping image patches for image classification and has achieved state-of-the-art accuracy. Since ViT models excel at capturing spatial information, they have also been extended to more challenging tasks, including action detection (Li et al, 2023), image retrieval (Xiang et al, 2020) and segmentation (You et al, 2023). For example, UCTransNet (Wang et al, 2022b) replaced the skip-connection with the channel transformer (CTrans) module. Karimi et. al. (Karimi et al, 2021) applied self-attention between neighboring image patches by modifying the MHSA mechanism of vision transformers. Despite the excellent performance in multiple image segmentation tasks. Vision Transformers suffer from the problem of computational overload, which has never been solved.
To address this problem, in this work, we make a first attempt to take advantage of the frequency domain information to boost the performance of medical image segmentation task. To the best of our knowledge, this is the first attempt to leverage the Multi-axis External Weights mechanism (MEW) that can simultaneously capture more comprehensive global and local information. We hope this work could shed some new light on potential tasks for medical image community to move forward to move forward.
@@ 3.2 Multi-axis External Weights Block
where $i=1,2,3$ corresponds to the first three branches. $W_{(I,J)}$ and $\mathcal{F}_{(I,J)}$ denote the learnable external weight and the 2D DFT for the corresponding axes. When $i=1$, $I$ and $J$ represent Height-Width axes. When $i=2$, $I$ and $J$ present Channel-Width axes. When i = 3, $I$ and $J$ represent Channel-Height axes. $\odot$ is the element-wise product. ($\mathcal{F}^{-1}_{(I,J)}$ refers to the 2D inverse DFT. $Split$ and $Concat$ represent the split and concatenation operation along the channel dimension.
@@
As shown in Figure 2 (c), we replace SA in ViT with our MEW and obtain our MEWB, which can be expressed as Eq. 5 and Eq. 6.
@@
where, $GroupNorm$ is the group normalization (Wu and He, 2018) and $FFN$ denotes to the feed-forward layer.
@@ 4.2 Implementation details
Following previous works (Ruan et al, 2022; Wang et al, 2022c), for the ISIC17 and ISIC18 datasets, all images are resized to 256 × 256. For the Synapse and ACDC dataset, all images are resized to 224 × 224. To avoid overfitting, data augmentation is performed, including random flip and random rotation. The loss function is BceDice loss. We set batch size equal to 8, and utilize CosineAnnealingLR (Loshchilov and Hutter, 2016) as the scheduler. Following training procedure in  Loshchilov and Hutter (2017), the initial learning rate (lr), maximum epochs (ep), and optimizer (opt) for the three datasets are itemized below:
@@
Dataset Model mIoU(%)normal-$\uparrow$ DSC(%)normal-$\uparrow$
@@
ISIC18 UNet (Ronneberger et al, 2015) 77.86 87.55
@@ 4.3 Comparison with State-of-the-Arts
Model DSC(%)normal-$\uparrow$ HD95(mm)normal-$\downarrow$
@@
Model DSC(%)normal-$\uparrow$ HD95(mm)normal-$\downarrow$
@@ 4.4 Ablation Studies
DW $W_{(H,W)}$ $W_{(C,W)}$ $W_{(C,H)}$ $Random$ mIoU(%)normal-$\uparrow$ DSC(%)normal-$\uparrow$
@@
The fundamental concept behind our MEW-UNet architecture lies in its multi-axis operations. To evaluate the effectiveness of this approach, we gradually increase the number of multi-axis operations and analyze the performance in an ablation study, as presented in Table 4. It can be observed that the adoption of DW convolution alone results in a decrease of almost 2% in the mIoU metric. Similarly, when solely utilizing single-axis frequency-domain transformation, the mIoU metric also decreases by 1%. The results indicate that our multi-axis operations can effectively capture comprehensive frequency domain information and global knowledge. Additionally, even when multi-axis operations are utilized, there is still a slight degradation in performance when leaving out the DW convolution in the fourth branch, which further emphasizes the importance of local information. Finally, we demonstrate the efficacy of our External Weights Generator by only randomly initializing learnable weights without IRB, and the degradation in performance can be clearly seen, which underscores the importance of using IRB for transforming learnable weights.
@@ 4.5 Visualization
To go even further, we also give some qualitative results of our proposed Multi-axis External Weights UNet named MEW-UNet. Figure 3 provides some visualized segmentation results by ITK-SNAP (Yushkevich et al, 2006). For instance, our MEW-UNet has on hole and missing boundary compared with MT-UNet on the Synapse dataset. Besides, compared to TransFuse (Zhang et al, 2021), our model has more clearly boundary on the ISIC2018 dataset on this challenging segmentation task.
@@ Acknowledgments
This work was partially supported by the National Natural Science Foundation of China under Grant No.62301315 and Startup Fund for Young Faculty at SJTU under Grant No.23X010501967. The authors would like to thank the anonymous reviewers for their valuable suggestions and constructive criticisms.
@@
    This work was partially supported by the National Natural Science Foundation of China under Grant No.62301315 and Startup Fund for Young Faculty at SJTU under Grant No.23X010501967.

### v2

@@ 1 Introduction
Medical image segmentation (MIS) has great practical value as it can assist relevant medical staff in locating the lesion area and improve the efficiency of clinical treatment, which has been widely studied, with state-of-the-art methods training convolutional neural networks in a supervised fashion, predicting a label map for a given input image (Dolz et al, 2018; Isensee et al, 2021; Kamnitsas et al, 2017). For a new segmentation problem, models are typically trained from scratch, requiring substantial design and tuning. Currently, although the CNN-based methods have achieved excellent performance in the field of medical image segmentation, they still cannot fully meet the strict requirements of medical applications for segmentation accuracy. Image segmentation is still a challenge task in medical image analysis area. In recent years, UNet (Ronneberger et al, 2015), an encoder-decoder model based on U-shape architecture, has been extensively employed for MIS due to its efficient feature extraction capability in handling multi-modal data. To be more specific, the encoder is used to gradually reduce the spatial resolution of the feature map and extract high-level semantic information from the image, while the decoder progressively restores the spatial resolution of the feature map and fuses low-level and high-level features to generate precise segmentation results, which enables UNet to quickly identify fine details in images, even accurately detecting tiny objects. As a result, many studies have been conducted based on the U-shape architecture, such as UNet++ (Zhou et al, 2018), which reduces the semantic gap between the encoder and decoder by incorporating dense connections. Additionally, Att-UNet (Oktay et al, 2018) introduces a gating mechanism to enable the model to focus on specific targets.
@@
The above improvements are all based on CNNs, which suffer from poor global information acquisition due to the natural locality of the convolution operation. In contrast, ViT (Dosovitskiy et al, 2020a) incorporates a self-attention mechanism (SA) that enhances the modeling ability of long-range dependencies and effectively captures holistic image semantic information, making it an ideal choice for intensive prediction tasks like image segmentation. As a result, recent advances can be categorized into two types. On the one hand, hybrid structures that combine CNNs and ViTs have gained popularity. For example, UCTransNet (Wang et al, 2022a) replaces the skip connection in UNet with the CTrans module, alleviating the issue of incompatible features between the encoder and decoder. MT-UNet (Wang et al, 2022c) employs CNNs at a shallow level and Local-Global SA along with the external attention mechanism at a deep level, to obtain richer representation information. On the other hand, some studies have leveraged pure ViTs, such as Swin-UNet (Cao et al, 2021), which replaces the convolution operation in U-Net with the Swin Transformer Block and achieves superior results. Although the aforementioned models focusing on spatial domain have achieved notable success in medical image area or other vision area (Ruan and Xiang, 2024; Xiang et al, 2022, 2023b, 2020b), the spatial domain modeling generally suffers from the low efficiency, e.g. it is very difficult to linearize the complexity of the vision Transformer from the perspective of spatial domain, which make them less favored by many image downstream tasks.
@@
In essence, the existing medical image segmentation methods mainly include spatial technique and frequency technique, among which spatial technique has some disadvantages. However, frequency domain technology has improved these problems. In general computer vision, feature extraction in the frequency domain has been shown to be a powerful approach (Zhong et al, 2022; Rao et al, 2021) to achieve more efficient and effective encoding and decoding processes. For instance, GFNet (Rao et al, 2021) leverages 2D discrete Fourier transform (2D DFT) to convert features from the spatial domain to the frequency domain, and filters in the frequency domain are used for learning representation. For MIS, many pathological areas are difficult to extract in the spatial domain, while they can be more easily extracted in the frequency domain (Huang et al, 2021). However, previous methods extract the frequency domain information only in a single axis, resulting in some signals that are still difficult to distinguish, which can be explained in Figure 1. In Figure 1 (a), it is challenging to visually distinguish the three regions from the spatial domain perspective. In contrast, Figure 1 (b) illustrates the signal strength of the three regions in the frequency domain, revealing relatively apparent differences in signal strength among them. However, there are still some signal strength intersections between the green and red curves in the case of employing DFT only for a single axis (the Height-Width axis). By considering the frequency domain signal strength of the three axes together, as shown in Figure 1 (c), it becomes evident that there is no signal strength intersection among the three curves. Motivated by this observation, we propose to extract and fuse features using a multi-axis approach.
@@
    The Multi-Axis External Weights block is proposed for the first time to obtain global and local information simultaneously, and frequency domain feature signals are introduced to comprehend the context more effectively.
@@
The remainder of this paper is structured as follows. In Section 2, we give the related works based on multi-axis representation learning in medical image segmentation, and then briefly introduce our method. In Section 3, The details of our multi-axis external weights UNet based on the U-shape architecture are presented. Extensive evaluations compared with state-of-the-art methods and comprehensive analyses of the proposed approach are reported in Section 4. Finally, conclusion of this paper and discussion of future works are presented in Section 5.
@@ 2.1 Medical Image Segmentation
First, a lot of endeavours have been concentrating on the segmentation process. These methods conquer different restrictions on conventional medical division techniques. Yet, there isn’t so much as one single strategy to be considered a better technique for different kinds of images; those techniques are only suitable for particular images and other applications (Xiang et al, 2023d; Gao et al, 2024).
@@ 2.2 ViT Based Techniques
Transformer models (Dosovitskiy et al, 2020b) entirely rely on the self-attention mechanism to build long-distance dependencies, which have achieved great success in almost all natural language processing tasks (Gao et al, 2023). Vision transformer (ViT) (Xiang et al, 2023a) is one of the earlier attempts to introduce transformer models into vision tasks, which applies a pure transformer architecture on non-overlapping image patches for image classification and has achieved state-of-the-art accuracy. Since ViT models excel at capturing spatial information, they have also been extended to more challenging tasks, including action detection (Li et al, 2023), image retrieval (Xiang et al, 2020a) and segmentation (You et al, 2023). For example, UCTransNet (Wang et al, 2022b) replaced the skip-connection with the channel transformer (CTrans) module. (Karimi et al, 2021) applied self-attention between neighboring image patches by modifying the MHSA mechanism of vision transformers. As for the model design, Duan et al (2023) proposed a hybrid transformer named DUCT consisting of dynamic local enhancement, unary co-occurrence excitation, and a standard multi-head self-attention module to learn the local, mid-level and global information, respectively. And a deep semi-supervised HAR approach named MixHAR (Duan et al, 2024) is also introduced to employ a linear interpolation mechanism to blend labelled and unlabelled activities while addressing both inter- and intra-activity variability. Despite the excellent performance in multiple image segmentation tasks. Vision Transformers suffer from the problem of computational overload, which has never been solved.
To address this problem, in this work, we make a first attempt to propose Multi-Axis External Weights block to obtain the frequency domain information on medical image. To the best of our knowledge, this is the first attempt to leverage the Multi-axis External Weights mechanism (MEW) that can simultaneously capture more comprehensive global and local information. We hope this work could shed some new light on potential tasks for medical image community to move forward to move forward.
@@ 3.2 Multi-axis External Weights Block
where $i=1,2,3$ corresponds to the first three branches. $W_{(I,J)}$ and $\mathcal{F}_{(I,J)}$ denote the learnable external weight and the 2D DFT for the corresponding axes. DW represents the depthwise separable convolution operation. When $i=1$, $I$ and $J$ represent Height-Width axes. When $i=2$, $I$ and $J$ present Channel-Width axes. When i = 3, $I$ and $J$ represent Channel-Height axes. $\odot$ is the element-wise product. ($\mathcal{F}^{-1}_{(I,J)}$ refers to the 2D inverse DFT. $Split$ and $Concat$ represent the split and concatenation operation along the channel dimension respectively.
@@
As shown in Figure 2 (c), we replace SA in ViT with our MEW and obtain our MEWB, which can be expressed as Eq. 5 and Eq. 6 respectively.
@@
where MEW denotes the multi-axis external weights block and $GroupNorm$ represents the group normalization (Wu and He, 2018). Besides, $FFN$ indicates the feed-forward layer. It should be noted that we adopt GroupNorm instead of LayerNorm in the vanilla ViT. This is due to the fact that we directly utilize the input $X\in\mathbb{R}^{C\times H\times W}$ instead of performing patch embedding operation. Furthermore, medical image datasets often have small sample sizes, and the batch size used in experiments is also typically small. In our network, we apply GroupNorm with four groups to mitigate the influence of batch size in BatchNorm (Ioffe and Szegedy, 2015).
@@ 4.2 Implementation details
Following previous works (Ruan et al, 2022; Wang et al, 2022c), for the ISIC17 and ISIC18 datasets, all images are resized to 256 × 256. For the Synapse and ACDC dataset, all images are resized to 224 × 224. To avoid overfitting, data augmentation is performed, including random flip and random rotation. The loss function is BceDice loss. We set batch size equal to 8, and utilize CosineAnnealingLR (Loshchilov and Hutter, 2016) as the scheduler. Following the parameter setting in Loshchilov and Hutter (2017), the initial learning rate (lr), maximum epochs (ep), and optimizer (opt) for the three datasets are itemized below:
@@
Dataset Model mIoU(%$\uparrow$ DSC(%$\uparrow$
@@
ISIC18 UNet Ronneberger et al (2015) 77.86 87.55
@@ 4.3 Comparison with State-of-the-Arts
Model DSC(%$\uparrow$ HD95(mm)$\downarrow$
@@
Model DSC(%$\uparrow$ HD95(mm)$\downarrow$
@@ 4.4 Ablation Studies
DW $W_{(H,W)}$ $W_{(C,W)}$ $W_{(C,H)}$ $Random$ mIoU(%$\uparrow$ DSC(%$\uparrow$
@@
Table 5: Ablation Studies for our MEWB on the ACDC dataset. $Random$ represents only randomly initialized learnable weights without $IRB$.
DW $W_{(H,W)}$ $W_{(C,W)}$ $W_{(C,H)}$ $Random$ DSC(%$\uparrow$ HD95(mm)$\downarrow$
90.05 1.65
90.22 1.30
90.44 1.25
90.54 1.22
91.00 1.19
90.23 1.29
The fundamental concept behind our MEW-UNet architecture lies in its multi-axis operations. To evaluate the effectiveness of this approach, we gradually increase the number of multi-axis operations and analyze the performance in an ablation study, as presented in Table 4. It can be observed that the adoption of DW convolution alone results in a decrease of almost 2% in the mIoU metric. Similarly, when solely utilizing single-axis frequency-domain transformation, the mIoU metric also decreases by 1%. The results indicate that our multi-axis operations can effectively capture comprehensive frequency domain information and global knowledge. Additionally, even when multi-axis operations are utilized, there is still a slight degradation in performance when leaving out the DW convolution in the fourth branch, which further emphasizes the importance of local information. On the other hand, the same phenomenon can also be observed when our Multi-axis External Weights block is tested on the ACDC datasets. As illustrated in Table 5, there also exists obvious performance degradation of 0.78% in terms of Dice Similarity Coefficient when only applied with the single-axis frequency-domain transformation. This confirms that frequency domain information is a crucial ingredient for increasing performance to an excellent level. Finally, we demonstrate the efficacy of our External Weights Generator by only randomly initializing learnable weights without IRB, and the degradation in performance can be clearly seen, which underscores the importance of using IRB for transforming learnable weights.
@@ 4.5 Visualization
To go even further, we also give some qualitative results of our proposed Multi-axis External Weights UNet named MEW-UNet. Figure 3 provides some visualized segmentation results by ITK-SNAP (Yushkevich et al, 2006). For instance, our MEW-UNet has on hole and missing boundary compared with MT-UNet on the Synapse dataset. Besides, compared to TransFuse (Zhang et al, 2021), our model has more clearly boundary on the ISIC2018 dataset on this challenging segmentation task. On the other hand, we also visualize the attention feature maps of the proposed Multi-axis External Weights Block (MEWB) on the ISIC2018 dataset, according to the Figure 4, we can easily observe that the attention maps with MEWB demonstrate a more precise focus on the lesion regions compared to the ones without MEWB. The attention overlay highlights how the model with MEWB captures the intricate details and boundaries of the lesions more effectively, leading to improved segmentation performance.
@@
Figure 4: The visualization comparison of the proposed Multi-axis External Weights Block on ISIC2018 dataset.
@@ Acknowledgments
This work was partially supported by the National Natural Science Foundation of China under Grant No.62301315, Startup Fund for Young Faculty at SJTU (SFYF at SJTU) under Grant No.23X010501967 and Shanghai Municipal Health Commission Health Industry Clinical Research Special Project under Grant No.202340010. The authors would like to thank the anonymous reviewers for their valuable suggestions and constructive criticisms.
@@
    This work was partially supported by the National Natural Science Foundation of China under Grant No.62301315, Startup Fund for Young Faculty at SJTU (SFYF at SJTU) under Grant No.23X010501967 and Shanghai Municipal Health Commission Health Industry Clinical Research Special Project under Grant No.202340010.